Write the expression as the​ sine, cosine, or tangent of a​ double-angle. Then find the exact value of the expression.
cos^2 135 degrees - sin^2 135 degrees

Answer by ikleyn(52794)

.
cos^2(135°) - sin^2(135°) = cos(135°)*cos(135°) - sin(135°)*sin(135°) =  (apply the adding formula for cosine) = 

= cos(135° + 135°) = cos(270°) = 0.

Solved.